OpenStudy (anonymous):

Calculate by hand the inverse of the matrix A = see my next post for the matrix A. Check whether your answer is correct. Use the inverse to solve both of the following sets of equations: x +z = 2 x + y +z = 3 2x + 3y+z = 5 y +z = 4 x + y +z = 5 3x + y+2z = 9 (Hint: rename the variables in the second system of equations.)

OpenStudy (anonymous):

\[A=\left[\begin{matrix}1 & 0 & 1 \\ 1 & 1 & 1 \\ 2 & 3 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

wow sorry x +z = 2 x + y +z = 3 2x + 3y+z = 5 there is suppose to be a gap here. y +z = 4 x + y +z = 5 3x + y+2z = 9

OpenStudy (amistre64):

cofactor; transpose, divide by det = inverse

OpenStudy (amistre64):

if the det = 0, you can prolly see why a matrix would not be invertible then

OpenStudy (anonymous):

yep know how to find the inverse...how do you do the second part

OpenStudy (anonymous):

oh they mean find the inverse of the next 2 system of equations...not the previous inverse

OpenStudy (anonymous):

not use the previous inverse

OpenStudy (anonymous):

i mean

OpenStudy (amistre64):

\[Ax=b\] \[x=A^{-1}b\]

OpenStudy (amistre64):

since youve found the inverse; they want you to define the vector for "x" i believe

OpenStudy (anonymous):

Inverse matrix can be calculated here step by step: http://www.emathhelp.net/calculators/inverse-of-matrix-calculator?i=%5B%5B1%2C0%2C1%5D%2C%5B1%2C1%2C1%5D%2C%5B2%2C3%2C1%5D%5D